# Experimental identification of topological defects in 2D colloidal glass

### v1

@@ Introduction
Figure 5: a-c Pair correlation functions for positive-positive, positive-negative and negative-negative defect pairs ($g_{\rm PP}(r)$, $g_{\rm PN}(r)$ and $g_{\rm NN}(r)$), respectively, for different $\omega$ values. Plots other than $\omega=50$ are shifted by the multiple of $0.5$ for clarity. d-e Spatial correlation between soft spots (with top 10% softness) and topological defects with d positive charge $g_{\rm PS}(r)$, and e negative charge $g_{\rm NS}(r)$, for different $\omega$ values. For better clarity, curves other than $\omega=15$ are shifted upward by the multiple of $0.25$. f Color map of charge density field obtained by the $\omega^{-1}$ weighted averaging of topological charge over frequencies in the range $500<\omega<$ 50, shown together with the soft spots (open symbols).
@@
To understand the spatial organization of defects, we calculate partial pair correlation functions $g_{\rm PP}(r)$, $g_{\rm PN}(r)$ and $g_{\rm NN}(r)$ for positive-positive, positive-negative and negative-negative defect pairs respectively, as shown in Fig. 5. From the positive-positive and negative-negative defect correlation in Fig. 5a and c, it is clear that, for small $r$, there is no correlation (or even a “depletion” zone near contact), indicating the absence of defects with similar charge in the vicinity of a tagged defect. Instead, the positive-negative correlation in Fig. 5b shows a very high probability for the presence of another defect with opposite charge in the neighborhood of a tagged defect. This observation suggests that defect charges with the same sign repel each other and those with opposite sign attract each other. It appears that the repulsion between two defects both with negative sign is slightly stronger because the repulsive “hard-core” distance at small separation $r$ in $g_{\rm NN}$ is bigger than that in $g_{\rm PP}$. Overall, both attraction and repulsion become weaker with the increase in frequency $\omega$, and at sufficiently high $\omega$ such defects are expected to be completely uncorrelated and homogeneously distributed.
@@
In recent studies, it has been possible to identify “soft spots” in the low-frequency vibrational modes where mesoscale relaxation or rearrangements are prone to happen [44, 78]. We identify such soft spots in the present system using the softness field defined in [78] (see Methods for details) and calculate their radial pair correlation with positive $(g_{\rm PS}(r))$ and negative $(g_{\rm NS}(r))$ defects (see Methods and SI). Figures 5d-e show the correlation for $\omega=15,20,30,50,100$. It is quite evident that the defects with negative charge are highly correlated with soft spots at small $r$, while defects with positive charge show no such correlation. This means negative charge defects tend to appear in the close vicinity of soft spots. Furthermore, it will be expected that these defects with negative charge will most likely be associated with mobile regions in structural relaxation or plastic rearrangements under deformation. Defects with positive charge located somewhat further away from the soft spots are also correlated, evidenced by the peak in $g_{\rm PS}$ at $r\approx 40$ $\mu$m. This is expected because we have already seen a significant correlation between +1 and -1 defects at the same length-scale, see Fig. 5b. With the increase in $\omega$, the correlation becomes weaker with both types of defects.
@@
Based on our experimental findings and analysis, we have successfully explored the topological characteristics of a two-dimensional colloidal glass system under the influence of an external magnetic field. Through meticulous control of experimental parameters and advanced numerical techniques, we have revealed intriguing insights into the interplay between topology, vibrational properties, and defect dynamics within disordered systems.
@@
By analyzing the structural properties of the topological defects, we have shown that defects of opposite charge pair together while defects with same charge repel each other, as expected from our intuition from electromagnetism. More importantly, we have observed a strong correlation at short distance between defects with negative topological charge and soft spots, confirming a close relation between anti-vortices and plasticity as suggested in [51]. Interestingly, we have found that a correlation between positive defects and plastic spots exist at larger length-scales as well, and it is possibly due to the positive correlation between negative and positive defects appearing at the same scale.
Our experimental analysis complements and confirms several of the results previously obtained from simulations at zero temperature [51], but it also provides several important new lessons and clarifications. First, it proves that the topological vortex-like defects proposed in [51] survive at finite temperature. Second, it clarifies that the number of these defects in realistic, and necessarily finite temperature, systems does not scale quadratically with frequency, as argued in [51], but rather linearly with it, nicely correlating with the vibrational properties of the system. Finally, our experimental data supports the claim of [51] of a strong local correlation between the -1 defects and plastic spots at short scales. Nevertheless, they also reveal a strong correlation with the +1 defects at an intermediate larger scale, that is consistent, and indeed expected, from the strong correlation among +1 and -1 defects. This suggests that +1 and -1 defects tend to pair together and create dipole structures, with possible intriguing connections with recent results about plastic screening in amorphous systems by Lemaître et al. [57].
@@ Methods
For each eigenvector field $(e^{x}_{i},e^{y}_{i})$ at $\omega_{i}$ ($i=1,2,...,2N$), we assign an angle $\theta(\vec{r})$ on every site at $\vec{r}$ of a $54\times 40$ rectangular lattice having grid length $r_{c}$ in both the directions and superposed to the experimental system. We obtain the phase angle $\theta(\vec{r})$ at each lattice site as
@@
where $\vec{d\ell}$ represents the line element along the closed square loop. Typically, the value of $q$ is an integer, and it is used to identify the locations of defects, which are characterized by non-zero values of $q$ at the center of the smallest square regions. For the calculation of topological charge, we do not enforce the periodic boundary conditions. We calculate the total number of defects $N_{d}$ by counting the defects with $q=+1$ and $-1$.
@@
The Pearson correlation coefficient ($C(\omega)$) is calculated as
@@
We calculate the softness field $\phi_{i}$ defined for each particle $i$ by superposing the participation fraction weighted by the corresponding mode energy in the low-frequency vibrational modes,
$\phi_{i}=\frac{1}{N_{m}}\sum_{j=1}^{N_{m}}\frac{|e^{(i)}_{j}|^{2}}{m_{i}\omega_{j}^{2}}.$ (7)
Here $e^{(i)}_{j}$ is the components corresponding to particle $i$ of the eigenvector associated with frequency $\omega_{j}$ and $N_{m}$ is the number of low energy modes considered for calculation. In this study, $N_{m}$ is considered such that $500<\omega<$ 50. Further, we identify the particles with top $10\%$ softness value as soft spots. The radial pair correlation function between the soft spots and the topological defects with positive charge in mode $l$ is defined as
@@
The code that support the findings of this study is available upon reasonable request by contacting the corresponding authors.
@@ Acknowledgements
We would like to thank Yuliang Jin, Deng Pan, Piotr Surowka, Jie Zhang, Cunyuan Jiang, Yujie Wang, Zihan Zheng, Jack Douglas, Tim W. Sirk and Chengran Du for many discussions on topological defects in glasses and related collaborations. M.B. acknowledges the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01) and the sponsorship from the Yangyang Development Fund. A.C.Y.L. acknowledges the support of the Australian Research Council (FT180100594). A.Z. and V.V. gratefully acknowledge funding from the European Union through Horizon Europe ERC Grant number: 101043968 “Multimech”. A.Z. gratefully acknowledges the Niedersächsische Akademie der Wissenschaften zu Göttingen in the frame of the Gauss Professorship program. A.Z. and A.B. gratefully acknowledge funding from US Army Research Office through contract nr. W911NF-22-2-0256.
@@
P. K. provided the dataset of the experiment. V. V and A. B. performed the numerical analysis. A. Z. and M. B supervised the project. All authors contributed to the writing of the manuscript.
@@
The authors declare that no competing interests exist.
Supplementary Information
In this Supplementary Information (SI) we provide additional details about the experimental conditions and setup, a brief description of the structure and dynamics of the system, and some simulation results to complement the experimental observation.
Additional information related to the experiment
Experimental information has been described in detail in [64], but we briefly outline here the crucial steps. Our system consists of a binary colloidal mixture having paramagnetic properties, suspended at a flat air-water interface due to gravity. The tilt and interface position is well controlled throughout the simulation to realize a perfect two-dimensional system. In the presence of an external magnetic field, for the typical inter-particle separation at room temperature, the dominating interaction between particles is dipole-dipole interaction.
Superparamagnetic colloidal particles: The colloidal particles are polystyrene beads doped with magnetite $({\rm Fe_{2}O_{3}})$ and sealed with a thin layer of epoxy. In the absence of an external magnetic field, net magnetization in each colloidal particle is zero. There are two types of particles (diameters: $d_{A}=4.5\pm 0.05\leavevmode\nobreak\ \mu{\rm m},d_{B}=2.8\leavevmode\nobreak\ \mu{\rm m}$; mass density: $\rho_{A}=1.5\leavevmode\nobreak\ {\rm g/cm^{3}},\rho_{B}=1.3\leavevmode\nobreak\ {\rm g/cm^{3}}$ and magnetic susceptibility per particle: $\chi_{A}=6.22\times 10^{-11}\leavevmode\nobreak\ {\rm Am^{2}/T},\chi_{B}=6.6\times 10^{-12}\leavevmode\nobreak\ {\rm Am^{2}/T}$). Particles observed under transmission electron microscopy show that the magnetite is uniformly distributed over each particle volume and bigger particles are quite monodisperse. Smaller particles may have small polydispersity.
Suspensions of bigger and smaller particles are prepared separately and after treating properly to make sure that the suspension is uniform and there is no particle aggregation, they are mixed properly to get the mixture with the desired composition $\xi=N_{A}/(N_{A}+N_{B})$; $N_{A}$ and $N_{B}$ are the numbers of bigger and smaller species respectively. For the present colloidal system, we have an equal number of bigger and smaller species, i.e., $\xi\approx 0.5$.
Experimental setup: Formation of air-water interface and its control: A cylindrical hole of diameter 6 mm and depth 1 mm is created in a glass plate. A water droplet hangs inside this cylindrical geometry due to surface tension. The bottom air-water interface hosts the suspension of binary colloidal mixture after sedimentation due to gravity, while the top of the interface is properly sealed. A computer-controlled microsyringe attached horizontally controls the volume of the pended drop to ensure a flat interface. A schematic diagram of this setup is shown in Fig. 1 of the main text. The colloidal particles can freely diffuse over the interface, in contrast to other experimental setups where such a system is realized over a substrate, it is difficult to avoid the pinning of a few particles. The curvature and inclination of the interface have been adjusted over months to achieve a gradient-free density profile. There are copper coils that generate an almost uniform magnetic field, compensating the earth’s magnetic field, in the plane of the sample (deviation is less than 0.5% in the field of interest).
Video microscopy is employed to visualize the particles from the below side of the sample and an 8-bit charge-coupled device camera records the gray scale images of particles which are analyzed by a computer. There are approximately $10^{5}$ particles in the sample but the camera has the field of view of $\approx 1{\rm mm}^{2}\leavevmode\nobreak\ (1158\times 865\mu{\rm m}^{2})$ which can capture about 2300 particles. In this setup, the particle trajectories can be recorded for several days with the spatial resolution of $\approx 50\leavevmode\nobreak\ {\rm nm}$ and a time resolution of $\approx 1\leavevmode\nobreak\ {\rm s}$.
For the analysis presented in the main text, we have utilized $1000$ configurations sampled over a period of 2772 seconds. Throughout the experiments, room temperature is maintained.
Structure and dynamics of the colloidal system
As discussed in the main article, $\Gamma$, which is the ratio of magnetic dipole-dipole interaction energy to thermal energy $(k_{B}T)$:
$\Gamma=\frac{\mu_{0}}{4\pi}\frac{(\pi n)^{3/2}}{k_{B}T}[\xi\chi_{B}+(1-\xi)\chi_{A}]^{2}.$ (9)
Here, $n$ is the area density computed via Voronoi tesselation. $\Gamma$ for the sample used in this study is $423$. At this value of $\Gamma$, the structure and dynamics of the system are of a typical glass. We present below the pair-correlation function and mean-squared displacement (MSD) of the sample.
Partial pair-correlations in Fig. 6 indicate that the mixture is homogeneous and the structure is disordered. MSD plotted in Fig. 7 has an extended plateau, this means that the dynamics of the system is largely arrested. These features establish the typical glassy behavior of the samples used in this study.
Figure 6: Partial pair correlation functions $g_{AA}(r),g_{AB}(r)$ and $g_{BB}(r)$ calculated for the samples used in the study.
Figure 7: Mean squared displacement (MSD) as a function of time.
Charge density and Softness map
In order to calculate the topological charge density, we have divided the entire experimental box into square cells of size $a_{s}=20.78{\mu}m$ and then count the number of topological defects $np$ and $nn$ with winding numbers $+1$ and $-1$, respectively, for each $\omega\in(0,50)$. Then the average charge density inside each square cell is defined as
$\bar{c}=\frac{\sum_{i}(np-nn)/{\omega}_{i}}{\sum_{i}1/{\omega}_{i}}.$ (10)
Here we have used the weighted factor $1/{\omega}$ due to the linear dependence of the total number of defects at lower frequencies.
Additional simulation results
We have also performed a set of simulations of a two dimensional Lennard-Jones glass as that studied by Wu et al. in [51]. We have extended their analysis of vibrational spectra and topological features to finite temperature systems, to better compare with our experimental setup. We aim to understand whether the low-frequency behavior of $N_{d}$, i.e., the total number of topological defects changes for finite temperature glasses. We perform additional simulations of thermal glass with the athermal one using LAMMPS [79]. Thermalized glassy configurations are prepared by cooling the high-temperature glass at $T_{i}=5.0$ to the target temperatures $T=0.1,0.2,0.4$ over $10^{7}$ time step $(\delta t=0.001)$ and then annealed for further $10^{6}$ number of time steps. For the preparation of athermal configurations, we have followed the same protocol as discussed in [51]; energy of $T=0.1$ glassy configuration is minimized using conjugate gradient algorithm.
We calculate the vibrational density of states using the instantaneous configurations at different temperatures and characterize the topological defects in the eigenspace. The fraction of unstable modes grows with temperature, in contrast to the athermal case where such modes are absent. Our focus is on the small omega behavior of $N_{d}$ with the temperature change. In Fig. 8 we observe that, unlike the athermal case where $N_{d}\sim\omega^{\alpha}$ with $\alpha=2$, the scaling of the number of defects with frequency becomes linear at finite temperature. Our experimental colloidal system is a finite temperature glass, as the presence of imaginary modes in the vibrational spectrum confirms. Consistent with this simulation analysis, the scaling of the number of defects in the experimental system shows an exponent $\alpha$ close to $1$.
Figure 8: Vibrational density of states $D(\omega)$ and number of topological defects $N_{d}$ as a function of frequency $\omega$ at different temperatures in a two dimensional Lennard-Jones glass.

### v2

@@ Introduction
Figure 6: Structure of defect pairs and the spatial correlation between defects and soft spots. a-c, Pair correlation functions for positive-positive, positive-negative and negative-negative defect pairs ($g_{\rm PP}(r)$, $g_{\rm PN}(r)$ and $g_{\rm NN}(r)$), respectively, for different $\omega$ values. Plots other than $\omega=50$ are shifted by the multiple of $0.5$ for clarity. d,e, Spatial correlation between soft spots (with top 10% softness) and topological defects with d positive charge $g_{\rm PS}(r)$, and e negative charge $g_{\rm NS}(r)$, for different $\omega$ values. For better clarity, curves other than $\omega=15$ are shifted upward by the multiple of $0.25$. f, Color map of charge density field obtained by the $\omega^{-1}$ weighted averaging of topological charge over frequencies in the range $500<\omega<$ 50, shown together with the soft spots (open symbols).
@@
To understand the spatial organization of defects, we calculate partial pair correlation functions $g_{\rm PP}(r)$, $g_{\rm PN}(r)$ and $g_{\rm NN}(r)$ for positive-positive, positive-negative and negative-negative defect pairs respectively, as shown in Fig. 6. From the positive-positive and negative-negative defect correlation in Fig. 6a and c, it is clear that, for small $r$, there is no correlation (or even a “depletion” zone near contact), indicating the absence of defects with similar charge in the vicinity of a tagged defect. Instead, the positive-negative correlation in Fig. 6b shows a very high probability for the presence of another defect with an opposite charge in the neighborhood of a tagged defect. This observation suggests that defect charges with the same sign repel each other and those with opposite sign attract each other. It appears that the repulsion between two defects both with negative sign is slightly stronger because the repulsive “hard-core” distance at small separation $r$ in $g_{\rm NN}$ is bigger than that in $g_{\rm PP}$. Overall, both attraction and repulsion become weaker with the increase in frequency $\omega$, and at sufficiently high $\omega$ such defects are expected to be completely uncorrelated and homogeneously distributed.
@@
In recent studies, it has been possible to identify “soft spots” in the low-frequency vibrational modes where mesoscale relaxation or rearrangements are prone to happen [32, 60]. We identify such soft spots in the present system using the softness field defined in [60] and calculate their radial pair correlation with positive $(g_{\rm PS}(r))$ and negative $(g_{\rm NS}(r))$ defects (see Supplementary Materials for details). Figs. 6d-e show the correlation for $\omega=15,20,30,50,100$. It is quite evident that the defects with negative charge are highly correlated with soft spots at small $r$, while defects with positive charge show no such correlation. This means negative charge defects tend to appear in the close vicinity of soft spots. Furthermore, it will be expected that these defects with negative charge will most likely be associated with mobile regions in structural relaxation or plastic rearrangements under deformation. Defects with positive charge located somewhat further away from the soft spots are also correlated, evidenced by the peak in $g_{\rm PS}$ at $r\approx 40$ $\mu$m. This is expected because we have already seen a significant correlation between +1 and -1 defects at the same length-scale, see Fig. 6b. With the increase in $\omega$, the correlation becomes weaker with both types of defects. Fig. 6f shows soft spots (open sysmbols) with the color map of weighted charge density field obtained by the averaging of topological charge in the frequency range $500<\omega<$ 50 (see Methods for the definition). This visual representation demonstrates the correlation between the zones with high negative charge density and soft spots.
@@
Based on our experimental findings and analysis, we have successfully explored the topological characteristics of a two-dimensional colloidal glass system under the influence of an external magnetic field. Through meticulous control of experimental parameters and numerical techniques, we have revealed intriguing insights into the interplay between topology, vibrational properties, and defect dynamics within disordered systems.
@@
By analyzing the structural properties of the topological defects, we have shown that defects of opposite charge pair together while defects with same charge repel each other, as expected from our intuition from electromagnetism. More importantly, we have observed a strong correlation at short distance between defects with negative topological charge and soft spots, confirming a close relation between anti-vortices and plasticity as suggested in [38]. Interestingly, we have found that a correlation between positive defects and plastic spots exists at larger length-scales as well.
Our experimental analysis complements and confirms several of the results previously obtained from simulations at zero temperature [38], but it also provides several important new lessons and clarifications. First, it proves that the topological vortex-like defects proposed in [38] survive at finite temperatures. Second, it clarifies that the number of these defects in realistic, and necessarily finite temperature, systems does not scale quadratically with frequency, as argued in [38], but rather linearly with it, nicely correlating with the vibrational properties of the system. Finally, our experimental data supports the claim of [38] of a strong local correlation between the -1 defects and plastic spots at short scales. Nevertheless, they also reveal a strong correlation with the +1 defects at an intermediate larger scale, that is consistent, and indeed expected, from the strong correlation among +1 and -1 defects. This suggests that +1 and -1 defects tend to pair together and create dipole structures, with possible intriguing connections with recent results about plastic screening in amorphous systems by Lemaître et al. [42].
@@ Methods
For each eigenvector field $(e^{x}_{i},e^{y}_{i})$ at $\omega_{i}$ ($i=1,2,...,2N$), we assign an angle $\theta(\vec{r})$ on every site at $\vec{r}$ of a $54\times 40$ rectangular lattice having grid length $r_{c}$ in both the directions and superposed to the experimental system. We obtain the phase angle $\theta(\vec{r})$ at each lattice site as [38]
@@
where $\vec{d\ell}$ represents the line element along the closed square loop. Typically, the value of $q$ is an integer, and it is used to identify the locations of defects, which are characterized by non-zero values of $q$ at the center of the smallest square regions. For the calculation of topological charge, we do not enforce the periodic boundary conditions. We also show a low-frequency configuration at $\omega=1.53$ Hz for three different interpolation grid lengths ($r_{c}$) in Supplementary Fig. 5, to show the robustness of the defect identification against the choice of grid lengths [61]. Further, these configurations are presented using Schlieren patterns [62, 63], where defect locations are marked by the merging of distinct color intensities, while uniform color regions indicate defect-free areas (see Supplementary Fig. 5d–f). Such patterns are consistent with the visualization of defects identified using eq.5.
We calculate the total number of defects $N_{d}$ by counting the defects with $q=+1$ and $-1$. This defect behavior remains robust against slight variations in the interpolation grid length $r_{c}$ (see Supplementary Fig. 6). For a given mode of eigenfrequency $\omega$, the radial pair correlation between topological defects $\alpha$ and $\beta$ with $\{\alpha,\beta\}\in\{\rm P,N\}$, is defined as
$g_{\rm\alpha\beta}(r)=\frac{L_{x}L_{y}}{2\pi rN_{\alpha}N_{\beta}}\sum_{i=1}^{N_{\alpha}}\sum_{j=1}^{N_{\beta}}\delta(r-|\vec{r_{ij}}|).$ (6)
Here, $L_{x},L_{y}$ are the dimensions of the region in which correlation is calculated, $\{\rm P,N\}$ are the positive and negative topological defects, $N_{\alpha}$ and $N_{\beta}$ are the respective numbers of such defects, and $|\vec{r_{ij}}|$ is the distance between defect $i$ and $j$.
@@
The Pearson correlation coefficient ($C(\omega)$) is calculated as [64]
@@
We calculate the softness field $\phi_{i}$ for each particle $i$ by superposing the participation fraction weighted by the corresponding mode energy in the low-frequency vibrational modes [32, 60],
$\phi_{i}=\frac{1}{N_{m}}\sum_{j=1}^{N_{m}}\frac{|{\bf e}^{(i)}_{j}|^{2}}{m_{i}\omega_{j}^{2}}.$ (8)
Here ${\bf e}^{(i)}_{j}$ is the eigenvector corresponding to particle $i$ associated with frequency $\omega_{j}\in(0,50)$ and $N_{m}$ is the number of low energy modes within this frequency range. Further, we identify the particles with top $10\%$ softness value as soft spots.
The radial pair correlation function between the soft spots and the topological defects with positive charge in mode $l$ is defined as
@@
Calculation of charge density
To calculate the topological charge density, we divided the experimental box into square cells of size $a_{s}=20.78{\mu}m$. Within each cell, we counted the number of topological defects $n_{p}$ and $n_{n}$ with winding numbers $+1$ and $-1$, respectively, for $\omega_{i}\in(0,50)$. The average charge density within each square cell is defined as
$\bar{c}=\frac{\sum_{i}(n_{p}-n_{n})/{\omega}_{i}}{\sum_{i}1/{\omega}_{i}}.$ (10)
Here, we employed the weighting factor $1/{\omega}_{i}$ to account for the linear dependence of the total number of defects on lower frequencies.
@@
The codes that support the findings of this study is available upon reasonable request by contacting the corresponding authors.
@@ Acknowledgements
We would like to thank Yuliang Jin, Deng Pan, Piotr Surowka, Jie Zhang, Cunyuan Jiang, Yujie Wang, Zihan Zheng, Jack Douglas, Tim W. Sirk and Chengran Du for many discussions on topological defects in glasses and related collaborations. M.B. acknowledges the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01) and the sponsorship from the Yangyang Development Fund. A.C.Y.L. acknowledges the support of the Australian Research Council (FT180100594). P.K. acknowledges funding of the German Research Foundation within the Heisenberg-Program, project number 453041792. A.Z. and V.V. gratefully acknowledge funding from the European Union through Horizon Europe ERC Grant number: 101043968 “Multimech”. A.Z. gratefully acknowledges the Niedersächsische Akademie der Wissenschaften zu Göttingen in the frame of the Gauss Professorship program. A.Z. and A.B. gratefully acknowledge funding from US Army Research Office through contract nr. W911NF-22-2-0256.
@@
P.K. provided the dataset of the experiment. V.V. and A.B. performed the numerical analysis. A.Z. and M.B. supervised the project. All authors contributed to the writing of the manuscript.
@@
The authors declare no competing interests.